# Generalized Frobenius Manifold Structures on the Orbit Spaces of Affine Weyl Groups I

### v1

@@ 3.1. The orbit space and pencil generators
Suppose $\{z^{1},\dots,z^{\ell}\}$ be a set of proper generators of $\mathscr{A}^{W}$. We say that it is a set of pencil generators of $\mathscr{A}^{W}$ if the associated metric $g_{\lambda}=\bigl{(}g^{ij}_{\lambda}\bigr{)}$ depends linearly on $\lambda$, i.e., $g^{ij}_{\lambda}$ can be represented in the form
@@
where $\eta=(\eta^{ij})$ is non-degenerate at generic points of $\mathcal{M}$, and the contravariant components $\Gamma^{ij}_{g_{\lambda},k}=-g_{\lambda}^{ir}\Gamma_{g_{\lambda},rk}^{j}$ of the Levi-Civita connection of $g_{\lambda}$ can be represented in terms of that of $g$ and $\eta$ as follows:
@@
From the above definition we know that a set of pencil generators of $\mathscr{A}^{W}$ yields a flat pencil[10] of metrics $g,\eta$. Note that the flat metric $g$ is independent of the choice of proper generators, since
@@
We will present examples of pencil generators in Sect. 5 and in a subsequent paper [20]. Unless otherwise specified, in what follows we always assume that $\{z^{1},\dots,z^{\ell}\}$ is a set of pencil generators of $\mathscr{A}^{W}$. We will end this section by proving that we can construct a generalized Frobenius manifold structure on $\mathcal{M}$ with the given flat pencil of metrics.
3.2. Homogeneous flat coordinates
We are to show in this subsection the existence of a system of homogeneous flat coordinates for the flat metric $\eta$.
@@ 3.2. Homogeneous flat coordinates
Here $\Gamma_{\eta,ij}^{k}$ are the Christoffel symbols of $\eta$’s Levi-Cività connection, they are given by
@@
From Definition 2.3 we know that
@@
We are to use the above homogeneity property to prove the existence of homogeneous flat coordinates of $\eta$. To this end, let us first clarify the solution space of the system of equations (3.9) that we are considering. Let $\widehat{\mathcal{M}}$ be the universal covering of $\mathcal{M}\setminus D$ with
$D=\{z\in\mathcal{M}~{}\big{|}~{}\det\eta(z)=0\}.$ (3.13)
@@
$\mathscr{Q}(\widehat{\mathcal{M}})=\mathscr{O}(\widehat{\mathcal{M}})/\mathbb{C},$
where $\mathscr{O}(\widehat{\mathcal{M}})$ refers to the holomorphic function ring on $\mathcal{M}$. Denote the solution space of (3.9) by
$\mathcal{K}=\left\{t\in\mathscr{Q}(\widehat{\mathcal{M}})\mid\frac{\partial^{2}t}{\partial z^{i}\partial z^{j}}=\Gamma_{\eta,ij}^{k}\frac{\partial t}{\partial z^{k}},\ i,j=1,\dots,\ell\right\}.$
@@
Then for any homogeneous function $f$ of degree $n$, the following identity holds true:
@@
here $\mathcal{L}_{E}$ is the Lie derivative along $E$. Hence homogeneous functions can be seen as eigenvectors of $\mathcal{L}_{E}$, of which the eigenvalues are their degrees. Now we consider the action of $\mathcal{L}_{E}$ on the solution space $\mathcal{K}$.
@@
In what follows, we always assume that $\mathcal{L}_{E}|_{\mathcal{K}}$ is diagonalizable. This implies that $\eta$ admits $\ell$ homogeneous flat coordinates which we denote by $t^{1},\dots,t^{\ell}$. Let
@@
it follows that if $d_{\alpha}+d_{\beta}\not=\kappa$ then $\eta^{\alpha\beta}$ must vanish. Thus by performing, if necessary, a degree preserving linear transformation, we can choose the flat coordinates $t^{1},\dots,t^{n}$ of $\eta$ such that
@@
:$*\colon\{1,\dots,l\}\to\{1,\dots,l\},\quad\alpha\to\alpha^{*},$
@@
We note that $d_{1},\dots,d_{\ell}$ are the eigenvalues of the operator $\nabla^{\eta}E$ corresponding to the eigenvectors $\frac{\partial}{\partial t^{1}},\dots,\frac{\partial}{\partial t^{\ell}}$, so they are also the eigenvalues of the representation matrix $A$ of $\nabla^{\eta}E$ in the basis $\frac{\partial}{\partial z^{1}},\dots,\frac{\partial}{\partial z^{\ell}}$. Since
@@
Remark 3.1.
When we choose $\omega$ to be a fundamental weight, we conjecture that all $d_{\alpha}$ are positive numbers. In what follows, we will continue our construction by assuming this fact.
@@ 3.3. The Frobenius algebra
Denote by $\Gamma_{\alpha\gamma}^{\beta}$ the Christoffel symbols of the Levi-Civita connection of $g$ in the flat coordinates $t^{1},\dots,t^{\ell}$ of $\eta$, and by
$\Gamma_{\gamma}^{\alpha\beta}=-g^{\alpha\epsilon}\Gamma_{\epsilon\gamma}^{\beta}$ (3.25)
the contravariant components of the Levi-Civita connection of $g$. Then from (3.7) we know that $\Gamma_{\gamma}^{\alpha\beta}$ are also the contravariant components of the Levi-Civita connection of $g+\lambda\eta$ for each $\lambda\in\mathbb{C}$.
We define an algebra structure on $T^{*}\mathcal{M}$ by
@@
[…]
$\omega_{e}=-\sum_{r=1}^{\ell}\frac{1}{\kappa}m_{r}\mathrm{d}\log z^{r}.$ (3.27)
@@
The 1-form (3.27) is a left unity of the algebra structure defined by (3.26) on $T^{*}\mathcal{M}$.
@@
From the definition of the multiplication (3.26) it follows that
@@
$g^{\sharp}\colon T^{*}\mathcal{M}\to T^{**}\mathcal{M}=T\mathcal{M},\quad\omega\mapsto g(\omega,\cdot).$ (3.29)
@@
Note that the null-transformation $\varphi_{0}$ transforms the metric $a$ defined in (3.2) to $g$, hence $x^{1},\dots,x^{\ell}$ are the flat coordinates of $g$. Since the relation between the coordinates $z^{1},\dots,z^{\ell}$ and $t^{1},\dots,t^{\ell}$ is independent of $\lambda$, in the following proof of the validity of (3.30) we can assume that $z^{i}$ and $t^{\alpha}$ are related with $x^{1},\dots,x^{\ell}$ via the null-transformation $\varphi_{0}$. From (2.43) we obtain
$\varphi_{0}^{*}\omega_{e}=-\frac{2}{\kappa}\pi i\mathrm{d}(\omega,\hat{\mathbf{x}}),$ (3.31)
it follows that
@@
$=-2\pi i\frac{1}{d_{\beta}}\sum_{r,s=1}^{\ell}m_{s}\frac{1}{4\pi^{2}}(\mathrm{d}x^{r},\mathrm{d}x^{s})\nabla^{a}_{\frac{\partial}{\partial x^{r}}}\left(\left.\frac{\partial t^{\beta}}{\partial x^{p}}\right|_{\lambda=0}\right)\mathrm{d}x^{p}$
@@
where $\deg f^{\beta}=\kappa+d_{\beta}$
@@
$\Gamma_{\alpha}^{\beta\gamma}=\frac{d_{\gamma}}{d_{\beta}+d_{\gamma}}\frac{\partial g^{\beta\gamma}}{\partial t^{\alpha}}.$ (3.38)
@@
By multiplying $e_{\alpha}$ on both sides of the second relation of (3.35) and by using (3.34) we obtain
$\sum_{\epsilon}d_{\epsilon}t^{\epsilon}\Gamma_{\epsilon}^{\beta\gamma}=d_{\gamma}g^{\beta\gamma},$ (3.39)
@@
$\sum_{\epsilon}d_{\epsilon}t^{\epsilon}\frac{\partial}{\partial t^{\epsilon}}\left(\eta^{\beta\nu}\frac{\partial f^{\gamma}}{\partial t^{\nu}}\right)=(d_{\gamma}+d_{\beta})\eta^{\beta\nu}\frac{\partial f^{\gamma}}{\partial t^{\nu}}=d_{\gamma}g^{\beta\gamma}.$ (3.40)
@@
$(d_{\gamma}+d_{\beta})\eta^{\beta\nu}\frac{\partial^{2}f^{\gamma}}{\partial t^{\nu}\partial t^{\alpha}}=d_{\gamma}\frac{\partial g^{\beta\gamma}}{\partial t^{\alpha}},$ (3.41)
@@
[…]
@@
The Frobenius algebra on $T^{*}\mathcal{M}$ induces a Frobenius algebra structure on $T(M\setminus D)$ via the isomorphism
$\eta^{\sharp}\colon T^{*}\mathcal{M}\to T^{**}\mathcal{M}=T\mathcal{M},\quad\omega\mapsto\eta(\omega,\cdot),$ (3.43)
where $D$ is define by (3.13). Explicitly, the Frobenius algebra structure $(\cdot,\langle\cdot\,,\cdot\rangle,e)$ on $T(M\setminus D)$ is defined by
$X\cdot Y=\eta^{\sharp}\left(\eta^{\flat}(X)\cdot\eta^{\flat}(Y)\right),\quad X,Y\in\rm{Vect}(\mathcal{M}),$ (3.44)
@@
$e=\eta^{\sharp}(\omega_{e}),$ (3.46)
here $\eta^{\flat}$ is the inverse map of $\eta^{\sharp}$. From (3.14), (3.20) and (3.27) it follows that
@@
3.4. The generalized Frobenius manifold structure on $\mathcal{M}\setminus D$
The notion of a generalized Frobenius manifold is a certain generalization of Dubrovin’s definition of Frobenius manifolds [9], it satisfies all axioms of the definition of a Frobenius manifold but the flatness of the unit vector field [22].
@@ 3.4. The generalized Frobenius manifold structure on $\mathcal{M}\setminus D$
[…]
@@
[…]
@@
    is symmetric w.r.t. $W,X,Y,Z\in\rm{Vect}(M)$.
@@
    There exist an Euler vector field $E$ on $M$ such that $\nabla\nabla E=0$, $\nabla E$ is diagonalizable and
    $\mathcal{L}_{E}(X\cdot Y)$ $=(\mathcal{L}_{E}X)\cdot Y+X\cdot(\mathcal{L}_{E}Y)+X\cdot Y.$ (3.48)
@@
    Here $\nabla$ is the Levi-Civita connection of the flat metric, and $d$ is a constant whic is called the charge of $M$.
Let us show that the Frobenius algebra on $T(\mathcal{M}\setminus D)$ that is defined in the last subsection yields a generalized Frobenius manifold structure on $\mathcal{M}\setminus D$ with the Euler vector field
$\tilde{E}=\frac{1}{\kappa}E=\sum_{\alpha=1}^{\ell}\frac{d_{\alpha}}{\kappa}\frac{\partial}{\partial t^{\alpha}}$ (3.50)
and with charge $d=1$. The validity of the axiom FM1 follows from the flatness of the metric $\eta$, and that of the axiom FM2 follows from the relation (3.42) and the second relation of (3.36). To check the quasi-homogeneity of the Frobenius manifold structure required by the axiom FM3, we need to verify equations (3.48) and (3.49) which can be represented in the flat coordinates $t^{1},\dots,t^{\ell}$ as follows:
@@
00 0 $=\left(1-\frac{d_{\alpha}}{\kappa}-\frac{d_{\beta}}{\kappa}\right)\eta_{\alpha\beta}.$ (3.52)
@@
and $(\eta_{\alpha\beta})=(\eta^{\alpha\beta})^{-1}$ with $\eta^{\alpha\beta}=\eta(\operatorname{d\!}t^{\alpha},\operatorname{d\!}t^{\beta})$. The validity of the identities follows from (3.20) and the fact that $\deg\Gamma^{\alpha\beta}_{\gamma}=d_{\alpha}+d_{\beta}-d_{\gamma}$. Thus we have the following theorem.
@@
Suppose $\{z^{1},\dots,z^{\ell}\}$ be a set of pencil generators associated with an irreducible reduced root system $R$ and a fixed weight $\omega$, then there exists a generalized Frobenius manifold structure of charge $d=1$ on $\mathcal{M}\setminus D$, of which the flat metric is given by $\eta$ and the multiplication is defined by (3.44). Moreover, the unit vector field $e$ and the Euler vector field $\tilde{E}$ of this generalized Frobenius are given by (3.46) and (3.50) respectively, and the intersection form coincides with $g$.
@@
We only need to show that the intersection form of the generalized Frobenius manifold $\mathcal{M}\setminus D$ coincides with $g$, i.e.,
@@
Note that the flat coordinates $t^{1},\dots,t^{\ell}$ as functions of $z^{1},\dots,z^{\ell}$ live on the universal covering $\widehat{\mathcal{M}\setminus D}$ of $\mathcal{M}\setminus D$, hence $\pi_{1}(\mathcal{M}\setminus D)$ acts on those functions. Denote the stable subgroup of this action by
$H=\{\gamma\in\pi_{1}(\mathcal{M}\setminus D)\mid\gamma^{*}t^{\alpha}(p)=t^{\alpha}(p),~{}\forall p\in\widehat{\mathcal{M}\setminus D}\},$ (3.57)
then $H$ corresponds to a covering space $\overline{\mathcal{M}}$ of $\mathcal{M}\setminus D$. Concretely, we may define an equivalence relation $\sim$ on $\widehat{\mathcal{M}\setminus D}$ by
Here $\pi\colon\widehat{\mathcal{M}\setminus D}\to\mathcal{M}\setminus D$ is the canonical projection. Then we have
¯$\overline{\mathcal{M}}=\widehat{\mathcal{M}\setminus D}/\sim.$ (3.59)
We are to show in Corollary 4.3 that $t^{1},\dots,t^{\ell}$ are coordinates of $\overline{\mathcal{M}}$. Therefore, the generalized Frobenius manifold structure that we constructed above is defined on $\overline{\mathcal{M}}$ instead of on $\mathcal{M}\setminus D$. If there is no ambiguity, we usually identify $\overline{\mathcal{M}}$ with $\mathcal{M}\setminus D$ since they are locally isomorphic. However, when we discuss the monodromy group of the generalized Frobenius manifold structure in the next section, it is necessary to view this generalized Frobenius manifold as $\overline{\mathcal{M}}$. We will prove in the next section that $z^{1},\dots,z^{\ell}$ are polynomials of $t^{1},\dots,t^{\ell}$, so we have the canonical projection
$\overline{\pi}:~{}\overline{\mathcal{M}}\to\mathcal{M}\setminus D.$
Since $\overline{\mathcal{M}}$ is constructed from the root system $R$, the weight $\omega$ and the pencil generators $z^{1},\dots,z^{\ell}$, we will also denote this generalized Frobenius manifold by
¯$\overline{\mathcal{M}}=\mathcal{M}(R,\omega,\{z^{1},\dots,z^{\ell}\}).$ (3.60)
If there is no confusion on the selection of the pencil generators, we simply write it as
¯$\overline{\mathcal{M}}=\mathcal{M}(R,\omega).$ (3.61)
4. Polynomial Property and the Monodromy Groups
In this section, we consider the monodromy group of the generalized Frobenius manifold ¯$\overline{\mathcal{M}}=\mathcal{M}(R,\omega)$. Note that the monodromy groups of the Frobenius manifold structures that are constructed in [9, 12] on the orbit spaces of the Coxeter groups and the extended affine Weyl groups coincide with the underlined groups. In the present case the monodromy group $\mathrm{Mono}(\overline{\mathcal{M}})$ can be embedded into the affine Weyl group $W_{a}(R)$. More precisely, we will show that the image of this embedding satisfies
$\mathrm{Mono}(\overline{\mathcal{M}})~{}\leq~{}\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell},$ (4.1)

### v3

@@ 3.1. The orbit space and pencil generators
Suppose $\{z^{1},\dots,z^{\ell}\}$ be a set of proper generators of $\mathscr{A}^{W}$. We say that it is a set of pencil generators of $\mathscr{A}^{W}$ if the associated metric $g_{\lambda}=\bigl(g^{ij}_{\lambda}\bigr)$ and its Christoffel symbols $\Gamma_{g_{\lambda},k}^{ij}$ depend at most linearly on $\lambda$, and the metric $\partial_{\lambda}g_{\lambda}^{ij}$ is non-degenerate at generic points.
From the above definition, for a set of proper generators $\{z^{1},\dots,z^{\ell}\}$, $g^{ij}_{\lambda}$ can be represented in the form
@@
where $\eta=(\eta^{ij})$ is non-degenerate at generic points of $\mathcal{M}$, and the Christoffel symbols $\Gamma^{ij}_{g_{\lambda},k}$ of $g_{\lambda}$ can be represented in terms of that of $g$ and $\eta$ as follows:
@@
Therefore, a set of pencil generators of $\mathscr{A}^{W}$ yields a flat pencil [10] of metrics $g,\eta$. Note that the flat metric $g$ is independent of the choice of proper generators, since
@@
We will present examples of pencil generators in Sect. 5 and in a subsequent paper [20]. Hereafter, $z^{1},\dots,z^{\ell}$} denotes a set of pencil generators for $\mathscr{A}^{W}$.
3.2. Quasi-homogeneous flat coordinates
We are to show in this subsection the existence of a system of quasi-homogeneous flat coordinates for the flat metric $\eta$.
@@ 3.2. Quasi-homogeneous flat coordinates
Here $\Gamma_{\eta,ij}^{k}$ are the Christoffel symbols of the Levi-Civita connection of $\eta$, they are given by
@@
From Definition 2.11 it follows that
@@
We are to use the above quasi-homogeneity property to prove the existence of quasi-homogeneous flat coordinates of $\eta$. To this end, let us first clarify the solution space of the system of equations (3.12). Let $\mathcal{W}$ be the universal covering space of $\mathcal{M}\setminus D_{\eta}$ with
$D_{\eta}=\{z\in\mathcal{M}\mid\det\eta(z)=0\}.$ (3.16)
@@
$\mathscr{Q}(\mathcal{W})=\mathscr{O}(\mathcal{W})/\mathbb{C},$
where $\mathscr{O}(\mathcal{W})$ refers to the holomorphic function ring on $\mathcal{W}$. Denote the solution space of (3.12) by
$\mathcal{K}=\left\{t\in\mathscr{Q}(\mathcal{W})\mid\frac{\partial^{2}t}{\partial z^{i}\partial z^{j}}=\Gamma_{\eta,ij}^{k}\frac{\partial t}{\partial z^{k}},\ i,j=1,\dots,\ell\right\}.$
@@
Then for any quasi-homogeneous function $f$ of degree $n$, the following identity holds true:
@@
here $\mathcal{L}_{E}$ is the Lie derivative along $E$. Hence quasi-homogeneous functions can be seen as eigenvectors of $\mathcal{L}_{E}$, of which the eigenvalues are their degrees. Now we consider the action of $\mathcal{L}_{E}$ on the solution space $\mathcal{K}$.
@@
In what follows, we assume for conciseness that $\mathcal{L}_{E}|_{\mathcal{K}}$ is diagonalizable. This assumption of diagonalization of the vector field $E$ implies that $\eta$ admits $\ell$ quasi-homogeneous flat coordinates which we denote by $t^{1},\dots,t^{\ell}$. Let
@@
it follows that $\eta^{\alpha\beta}$ vanishes if $d_{\alpha}+d_{\beta}\not=\kappa$. Thus by performing, if necessary, a degree preserving linear transformation, we can choose the flat coordinates $t^{1},\dots,t^{n}$ of $\eta$ such that
@@
$\colon\{1,\dots,l\}\to\{1,\dots,l\},\quad\alpha\to\alpha^{*},$
@@
We note that $d_{1},\dots,d_{\ell}$ are the eigenvalues of the operator $\nabla^{\eta}E$ on each point of $\mathcal{M}\setminus D_{\eta}$, corresponding to the eigenvectors $\frac{\partial}{\partial t^{1}},\dots,\frac{\partial}{\partial t^{\ell}}$, so they are also the eigenvalues of the representation matrix $A$ of $\nabla^{\eta}E$ in the basis $\frac{\partial}{\partial z^{1}},\dots,\frac{\partial}{\partial z^{\ell}}$. Since
@@
Lemma 3.5.
All $d_{\alpha}$ have positive real parts.
The proof of this lemma will be given in Appendix A. In fact, we will show in Corollary 4.4 that all $d_{\alpha}$ are positive rational numbers.
@@ 3.3. The Frobenius algebra
Denote by
$\Gamma_{\gamma}^{\alpha\beta}=-g^{\alpha\epsilon}\Gamma_{\epsilon\gamma}^{\beta},\quad\alpha,\beta,\gamma=1,\dots,\ell$ (3.28)
the Christoffel symbols of the Levi-Civita connection of the contravariant $g$ in the flat coordinates $t^{1},\dots,t^{\ell}$ of $\eta$. Then from (3.10) we know that $\Gamma_{\gamma}^{\alpha\beta}$ are also the Christoffel symbols of the Levi-Civita connection of $g+\lambda\eta$ for each $\lambda\in\mathbb{C}$. Moreover, we know from Remark 3.1 that $\Gamma^{\alpha\beta}_{\gamma}$ are defined globally on $\mathcal{M}\setminus D_{\eta}$.
We define an algebra structure on $T^{*}(\mathcal{M}\setminus D_{\eta})$ by
@@
where $d_{\beta}$ is the degree of $t^{\beta}$. From Lemma 3.5, we know in particular that $d_{\beta}\not=0$. This algebra is called the dual Frobenius algebra on $T^{*}(\mathcal{M}\setminus D_{\eta})$. Denote
$D_{0}=\{z\in\mathcal{M}\,|\,\exists\,i\in S,\,z^{i}=0\},\quad D=D_{\eta}\cup D_{0},$
and $\mathcal{M}_{D}=\mathcal{M}\setminus D$, then $\mathcal{M}_{D}$ is a dense open subset of $\mathcal{M}$. Let us define a one-form
$\omega_{e}=-\sum_{r=1}^{\ell}\frac{1}{\kappa}m_{r}\mathrm{d}\log z^{r}$ (3.30)
on $\mathcal{M}_{D}$. We are to show in this subsection that the multiplication (3.29) endows $T^{*}\mathcal{M}_{D}$ a Frobenius algebra structure with the unity $\omega_{e}$ and the bilinear form $\langle\cdot\,$$\cdot\rangle$ given by the flat metric $\eta$.
@@
The one-form (3.30) is a left unity of the algebra structure defined by (3.29) on $T^{*}\mathcal{M}_{D}$.
@@
As a closed condition, it suffices to prove this proposition on generic points of $\{\det g\not=0\}$ on $\mathcal{M}_{D}$. From the definition of the multiplication (3.29) it follows that
@@
$g^{\sharp}\colon T^{*}\mathcal{M}_{D}\to T^{**}\mathcal{M}_{D}=T\mathcal{M}_{D},\quad\omega\mapsto g(\omega,\cdot).$ (3.32)
@@
Since the relation between the coordinates $z^{1},\dots,z^{\ell}$ and $t^{1},\dots,t^{\ell}$ is independent of $\lambda$, in the following proof of the validity of (3.33) we can assume that $z^{i}$ and $t^{\alpha}$ are related with $x^{1},\dots,x^{\ell}$ via the null-transformation $\varphi_{0}$. Note that the null-transformation $\varphi_{0}$ transforms the constant metric $a$ to $g$, hence $x^{1},\dots,x^{\ell}$ are the flat coordinates of $g$. From (2.45) we obtain
[…]
so it follows that
@@
$=-2\pi i\frac{1}{d_{\beta}}\sum_{r,s=1}^{\ell}m_{s}\frac{1}{4\pi^{2}}a^{rs}\nabla^{a}_{\frac{\partial}{\partial x^{r}}}\left(\left.\frac{\partial t^{\beta}}{\partial x^{p}}\right|_{\lambda=0}\right)\mathrm{d}x^{p}$
@@
where $\deg f^{\beta}=\kappa+d_{\beta}$ .
@@
$\Gamma_{\gamma}^{\alpha\beta}=\frac{d_{\beta}}{d_{\alpha}+d_{\beta}}\frac{\partial g^{\alpha\beta}}{\partial t^{\gamma}}.$ (3.41)
@@
By multiplying $e_{\gamma}$ on both sides of the second relation of (3.38) and by using (3.37) we obtain
$\sum_{\epsilon}d_{\epsilon}t^{\epsilon}\Gamma_{\epsilon}^{\alpha\beta}=d_{\beta}g^{\alpha\beta},$ (3.42)
@@
$\sum_{\epsilon}d_{\epsilon}t^{\epsilon}\frac{\partial}{\partial t^{\epsilon}}\left(\eta^{\alpha\nu}\frac{\partial f^{\beta}}{\partial t^{\nu}}\right)=(d_{\beta}+d_{\alpha})\eta^{\alpha\nu}\frac{\partial f^{\beta}}{\partial t^{\nu}}=d_{\beta}g^{\alpha\beta}.$ (3.43)
@@
$(d_{\alpha}+d_{\beta})\eta^{\alpha\nu}\frac{\partial^{2}f^{\beta}}{\partial t^{\nu}\partial t^{\gamma}}=d_{\beta}\frac{\partial g^{\alpha\beta}}{\partial t^{\gamma}},$ (3.44)
@@
The algebra defined by the multiplication (3.29) on $T^{*}\mathcal{M}_{D}$ is a Frobenius algebra with the bilinear form $\langle\cdot\,$$\cdot\rangle$ given by the flat metric $\eta$ and with the unity (3.30).
@@
The Frobenius algebra on $T^{*}\mathcal{M}_{D}$ induces a Frobenius algebra structure on $T\mathcal{M}_{D}$ via the isomorphism
$\eta^{\sharp}\colon T^{*}\mathcal{M}_{D}\to T\mathcal{M}_{D},\quad\omega\mapsto\eta(\omega,\cdot),$ (3.46)
where $D$ is define by (3.16). Explicitly, the Frobenius algebra structure $(\cdot,\langle\cdot\,,\cdot\rangle,e)$ on $T\mathcal{M}_{D}$ is defined by
$X\cdot Y=\eta^{\sharp}\left(\eta^{\flat}(X)\cdot\eta^{\flat}(Y)\right),\quad X,Y\in\rm{Vect}(\mathcal{M}_{D}),$ (3.47)
@@
$e=\eta^{\sharp}(\omega_{e})=-\frac{1}{\kappa}\mathrm{grad}_{\eta}\sum_{r=1}^{\ell}m_{r}\log z^{r},$ (3.49)
where $\eta^{\flat}$ is the inverse map of $\eta^{\sharp}$. From (3.17), (3.23) and (3.30) it follows that
@@
3.4. The generalized Frobenius manifold structure on $\mathcal{M}_{D}$
A generalized Frobenius manifold extends Dubrovin’s original definition of a Frobenius manifold [9]; it satisfies all axioms of the definition of a Frobenius manifold but the flatness of the unit vector field [22].
@@ 3.4. The generalized Frobenius manifold structure on $\mathcal{M}_{D}$
A (complex) generalized Frobenius manifold (GFM) is a complex manifold $M$ endowed on each of its tangent spaces $T_{p}M$ a Frobenius algebra $A_{p}=(\cdot,\langle\cdot\,,\cdot\rangle,e)$ which depends holomorphically on $p$, here $\cdot$ is the operation of multiplication of the algebra, $\langle\cdot\,$$\cdot\rangle$ is a symmetric non-degenerate bilinear form which is compatible with the operation of multiplication, and $e$ is the unity. It is required to satisfy the following axioms:
@@
    The bilinear form $\langle\cdot\,$$\cdot\rangle$ yields a flat metric on $M$.
@@
    is symmetric w.r.t. $W,X,Y,Z\in\rm{Vect}(M)$, where $\nabla$ is the Levi-Civita connection of the flat metric.
@@
    There exists an Euler vector field $E$ on $M$ such that $\nabla\nabla E=0$, $\nabla E$ is diagonalizable and
    $\mathcal{L}_{E}(X\cdot Y)$ $=(\mathcal{L}_{E}X)\cdot Y+X\cdot(\mathcal{L}_{E}Y)+X\cdot Y,$ (3.51)
@@
    here $d$ is a certain constant called the charge of $M$.
Lemma 3.12.
The Frobenius algebra on $T\mathcal{M}_{D}$ defined in the last subsection yields a generalized Frobenius manifold structure on $\mathcal{M}_{D}$ with the Euler vector field
$\tilde{E}=\frac{1}{\kappa}E=\sum_{\alpha=1}^{\ell}\frac{d_{\alpha}t^{\alpha}}{\kappa}\frac{\partial}{\partial t^{\alpha}}$ (3.53)
and with charge $d=1$.
Proof.
The validity of the axiom FM1 follows from the flatness of the metric $\eta$, and that of the axiom FM2 follows from the relation (3.45) and the second relation of (3.39). To check the quasi-homogeneity of the Frobenius manifold structure required by the axiom FM3, we need to verify equations (3.51) and (3.52) which can be represented in the flat coordinates $t^{1},\dots,t^{\ell}$ as follows:
@@
$0$ $=\left(1-\frac{d_{\alpha}}{\kappa}-\frac{d_{\beta}}{\kappa}\right)\eta_{\alpha\beta}.$ (3.55)
@@
and $(\eta_{\alpha\beta})=(\eta^{\alpha\beta})^{-1}$ with $\eta^{\alpha\beta}=\eta(\operatorname{d\!}t^{\alpha},\operatorname{d\!}t^{\beta})$. The validity of the identities follows from (3.23) and the fact that $\deg\Gamma^{\alpha\beta}_{\gamma}=d_{\alpha}+d_{\beta}-d_{\gamma}$. ∎
Now we give the main theorem of this paper.
@@
Suppose $\{z^{1},\dots,z^{\ell}\}$ be a set of pencil generators associated with an irreducible reduced root system $R$ and a fixed weight $\omega$, and the vector field $E$ given by (3.17) is diagonalizable, then there exists a generalized Frobenius manifold structure of charge $d=1$ on $\mathcal{M}_{D}$, of which the flat metric is given by $\eta$ and the multiplication is defined by (3.47). Moreover, the unit vector field $e$ and the Euler vector field $\tilde{E}$ of this generalized Frobenius are given by (3.49) and (3.53) respectively, and the intersection form coincides with $g$.
@@
We only need to show that the intersection form of the generalized Frobenius manifold $\mathcal{M}_{D}$ coincides with $g$, i.e.,
@@
Since the generalized Frobenius manifold structure of $\mathcal{M}_{D}$ is constructed from the root system $R$, the weight $\omega$ and the pencil generators $z^{1},\dots,z^{\ell}$, we denote this generalized Frobenius manifold by
$\mathcal{M}_{D}=\mathcal{M}(R,\omega,\{z^{1},\dots,z^{\ell}\}).$
If there is no ambiguity in the selection of the pencil generators, we simply write it as
$\mathcal{M}_{D}=\mathcal{M}(R,\omega).$
4. Polynomiality Property and the Monodromy Groups
In this section, we consider the monodromy group of the generalized Frobenius manifold $\mathcal{M}_{D}=\mathcal{M}(R,\omega)$. Note that the monodromy groups of the Frobenius manifold structures that are constructed in [9, 12] on the orbit spaces of the Coxeter groups and the extended affine Weyl groups coincide with the underlined groups. In the present case the monodromy group $\mathrm{Mono}(\mathcal{M}_{D})$ can be embedded into the affine Weyl group $W_{a}(R)$. More precisely, we will show that the image of this embedding is
$\mathrm{Mono}(\mathcal{M}_{D})~\cong~\mathrm{Stab}_{W}(\omega)\ltimes\mathbb{Z}^{\ell},$ (4.1)